OpenStudy (anonymous):

Find the sum of the infinite series 1/3+1/9+1/27 if it exists.

OpenStudy (anonymous):

infinite? do you mean 1/3+1/9+1/27+1/81+...?

mathslover (mathslover):

@sara_singh12 can you tell me that is it GP or AP ?

mathslover (mathslover):

@Clouds yes I think it is \[\large{\frac{1}{3} + \frac{1}{9} + \frac{1}{27} + \frac{1}{81} + ...}\]

OpenStudy (anonymous):

in fact,\[p+p^2+p^3+...+p^n=\frac{ p-p^{n+1} }{ 1-p }\]

mathslover (mathslover):

I think @sara_singh12 is not present here , though I am putting the formula up here so that will help her to solve the problem : \[\large{S _ \infty = \frac{a}{1-r}}\]

OpenStudy (anonymous):

it is an infinite gp with common ratio of 1/3 so, sum = a/(1 - r) = 1/3 divided by ( 1 - 1/3) = 1/2

mathslover (mathslover):

More explanation First term = \(\frac{1}{3}\) Common ratio = \(\large{\frac{a_2}{a_1}}\) = \(\large{\frac{\frac{1}{9}}{\frac{1}{3}}}\) = \(\large{\frac{1}{3}}\) Therefore : \[\large{S_ \infty = \frac{\frac{1}{3}}{1- \frac{1}{3}}}\] \[\large{S_\infty = \frac{\frac{1}{3}}{\frac{2}{3}}}\] \[\large{S_\infty = \frac{1}{2}}\]